# GPU-accelerated Subcycling Time Integration with the Einstein Toolkit

### v1

@@ II Implementation Subcycling Within CarpetX
The values of ${\bf Y}_{i}$ are obtained from the Taylor expansion of ${\bf k}_{i}$:
@@
${\bf Y}_{3}$ $={\bf y}_{n}+\frac{h}{2}{\bf y}^{\prime}_{n}+\frac{h^{2}}{4}{\bf y}^{\prime\prime}_{n}+\frac{h^{3}}{16}({\bf y}^{\prime\prime\prime}_{n}-f_{y}{\bf y}^{\prime\prime}_{n})$ (17)
${\bf Y}_{4}$ $={\bf y}_{n}+h{\bf y}^{\prime}_{n}+\frac{h^{2}}{2}{\bf y}^{\prime\prime}_{n}+\frac{h^{3}}{8}({\bf y}^{\prime\prime\prime}_{n}+f_{y}{\bf y}^{\prime\prime}_{n}),$ (18)
where $f_{y}{\bf y}^{\prime\prime}_{n}=4({\bf k}_{3}-{\bf k}_{2})/h^{2}$. The key idea is to use dense output from the parent grid to evaluate ${\bf y}_{n}$, ${\bf y}^{\prime}_{n}$, ${\bf y}^{\prime\prime}_{n}$ and ${\bf y}^{\prime\prime\prime}_{n}$ in the AMR boundary ghost zones, and then apply Eqs. (15)-(18) to compute ${\bf Y}_{i}$ in the AMR ghost region.
@@
    Perform a full update on the coarse level using the physical boundary condition to compute $Y^{0}_{1}$, …, $Y^{0}_{4}$ at the boundary. Store $k^{0}_{1}$, …, $k^{0}_{4}$.
@@
    On the next finer level $\ell$ prolongate in space $k^{\ell-1}_{1}$, …, $k^{\ell-1}_{4}$ to mesh refinement boundary.
@@
    Perform a full update on level $\ell$. Use (10)-(18) with $\theta=0$ to compute $Y^{\ell}_{1}$, …, $Y^{\ell}_{4}$ at the mesh refinement boundary. Use (10) with $\theta=1/2$ to compute $y^{\ell}$ at mesh refinement boundary when implementing (5). Store $k^{\ell}_{1}$, …, $k^{\ell}_{4}$.
@@
    Perform a full update on level $\ell$. Use (10)-(18) with $\theta=1/2$ to compute $Y^{\ell}_{1}$, …, $Y^{\ell}_{4}$ at the mesh refinement boundary. Use (10) with $\theta=1$ to compute $y^{\ell}$ at mesh refinement boundary when implementing (5). Store $k^{\ell}_{1}$, …, $k^{\ell}_{4}$.
@@ III.1 Z4cow Thorn
To stabilize numerical evolution, we incorporate Kreiss-Oliger numerical dissipation [31] into the system. This is implemented by modifying the time derivative of the evolved variables as follows:
$\partial_{t}u\rightarrow\textsc{RHS}+Qu$ (35)
where $Qu$ represents the dissipation term, given by:
@@ IV.1 Scalar Wave
where $f(v)\equiv A\exp(-v^{2}/(2\sigma^{2}))$, with $A=1.0$, $\sigma=0.25/\sqrt{2}$. The scalar wave equation was evolved at three different resolutions: $\Delta x=h_{1},h_{1}/1.5,h_{1}/2$, where $h_{1}=1/8$. The computational domain utilized a three-level refinement grid, with refinement boundaries at $r=4.0$, $1.0$ and $0.5$, respectively.
@@
Figure 1: Errors in the solution of a Gaussian scaler wave. The plots show errors in low-, medium-, and high-resolution numerical solutions for both uniform time step and subcycling cases. The dashed lines represent the scaled medium- and high-resolution errors assuming fourth-order convergence, and they align closely with the low-resolution solution.
@@ IV.2.1 Convergence tests for Z4cow
Figure 2: Convergence test of a $q=1$, nonspinning binary using Z4cow. The top panel shows the low-, medium-, and high-resolution waveforms, which are visually indistinguishable, demonstrating excellent agreement across the resolutions. The middle and bottom panels display the rescaled differences between waveforms across various time periods, under the assumption of fourth-order convergence.
In Fig.2, the top panel illustrates the evolution of the real part of the $l=2$, $m=2$ mode of the Weyl scalar $\Psi_{4}$, extracted at a radius of $r=60M$, which encodes outgoing gravitational radiation. The evolution is shown over the time interval $t=300M$ to $t=600M$ for three different resolutions. Here, $\Psi_{4}$ is calculated using the WeylScal4 [37] and Multipole thorns in the SpacetimeX repository. At this scale, the results are visually indistinguishable, indicating excellent agreement across the resolutions.
The middle and bottom panels display the differences between the low- and high-resolution results, as well as the medium- and high-resolution results. Additionally, the difference between the medium- and high-resolution results is rescaled under the assumption of fourth-order convergence. Numerical oscillations are observed around $t=70M$, $t=200M$, and $t=300M$, which are attributed to numerical errors that do not follow fourth-order convergence (middle panel). However, once the gravitational wave signal becomes stronger and dominates, starting around $t=500$, the results exhibit clear fourth-order convergence (bottom panel). This behavior confirms the robustness and accuracy of the implementation in capturing the dominant physical features of the system.
@@ IV.2.2 Comparison with LazEv
Figure 3: A comparison of the Z4cow and LazEv waveforms. The top panel shows the amplitudes of the waveforms from the two codes. The next panel down shows the differences in the amplitudes between the two codes. The third panel show the waveform phases for the two codes. Finally, the bottom panel show the phase differences between the two codes.
@@ IV.2.3 Irreducible mass
Figure 4: Irreducible mass of the inspiraling binary black holes (top) and the merged black hole (bottom). As the resolution increases, mass conservation improves.
@@ IV.2.4 Constraint violations
Figure 5: Hamiltonian and Momentum constraints for three different resolution cases. The dashed lines represent the scaled medium- and high-resolution constraints, assuming fourth-order convergence. Fourth-order convergence is observed, except around $t=200M$ and $t=400M$.
@@ IV.3 Scaling Test
$224$ $3.31\times 10^{7}$ 1.00 $224$ $1.14\times 10^{8}$ 1.00 $4$ $1.82\times 10^{9}$ 1.00
$448$ $5.99\times 10^{7}$ 0.90 $448$ $1.96\times 10^{8}$ 0.86 $8$ $2.60\times 10^{9}$ 0.71
$896$ $9.07\times 10^{7}$ 0.69 $896$ $3.26\times 10^{8}$ 0.71 $16$ $3.48\times 10^{9}$ 0.48
$1792$ $1.12\times 10^{8}$ 0.42 $1792$ $5.40\times 10^{8}$ 0.59 $32$ $4.03\times 10^{9}$ 0.28
$3584$ $9.50\times 10^{7}$ 0.18 $3584$ $8.90\times 10^{8}$ 0.49 $64$ $3.97\times 10^{9}$ 0.14
@@
$224$ $8.26\times 10^{6}$ 1.00 $224$ $2.85\times 10^{7}$ 1.00 $4$ $4.55\times 10^{8}$ 1.00
$448$ $7.92\times 10^{6}$ 0.96 $448$ $2.64\times 10^{7}$ 0.92 $8$ $4.26\times 10^{8}$ 0.94
$896$ $7.94\times 10^{6}$ 0.96 $896$ $2.49\times 10^{7}$ 0.87 $16$ $3.96\times 10^{8}$ 0.87
$1792$ $7.32\times 10^{6}$ 0.89 $1792$ $2.48\times 10^{7}$ 0.87 $32$ $3.94\times 10^{8}$ 0.87
$3584$ $6.69\times 10^{6}$ 0.81 $3584$ $2.45\times 10^{7}$ 0.86 $64$ $3.70\times 10^{8}$ 0.81
@@ Acknowledgements.
The authors would like to thank Manuela Campanelli, Michail Chabanov, Jay V. Kalinani, Carlos Lousto and Weiqun Zhang for valuable discussions. The authors gratefully acknowledge the National Science Foundation (NSF) for financial support from grants OAC-2004044, OAC-2004157, OAC-2004879 PHY-2110338, OAC-2411068, PHY02409706, as well as the National Aeronautics and Space Administration (NASA) for financial support from TCAN Grant No. 80NSSC24K0100. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. RH acknowledges additional support by NSF grants OAC-2103680, OAC-2310548, and OAC-2005572. ES acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). ES remercie le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) de son soutien. This research used resources from the Texas Advanced Computing Center’s (TACC) Frontera and Vista supercomputer allocations (award PHY20010). Additional resources were provided by the BlueSky, Green Prairies, and Lagoon clusters of the Rochester Institute of Technology (RIT) acquired with NSF grants PHY-2018420, PHY-0722703, PHY-1229173, and PHY-1726215. We thank LSU HPC and CCT IT Services for making the mike compute resource available.

### v2

@@ II Implementation Subcycling Within CarpetX
The Taylor expansions of ${\bf Y}_{i}$ around ${\bf y}_{n}$ are given by:
@@
${\bf Y}_{3}$ $={\bf y}_{n}+\frac{h}{2}{\bf y}^{\prime}_{n}+\frac{h^{2}}{4}{\bf y}^{\prime\prime}_{n}+\frac{h^{3}}{16}({\bf y}^{\prime\prime\prime}_{n}-{\bf f}_{\bf y}{\bf y}^{\prime\prime}_{n})$ (18)
${\bf Y}_{4}$ $={\bf y}_{n}+h{\bf y}^{\prime}_{n}+\frac{h^{2}}{2}{\bf y}^{\prime\prime}_{n}+\frac{h^{3}}{8}({\bf y}^{\prime\prime\prime}_{n}+{\bf f}_{\bf y}{\bf y}^{\prime\prime}_{n}),$ (19)
where ${\bf f}_{\bf y}{\bf y}^{\prime\prime}_{n}=4({\bf k}_{3}-{\bf k}_{2})/h^{2}$. The key idea is to use dense output from the parent grid to evaluate ${\bf y}_{n}$, ${\bf y}^{\prime}_{n}$, ${\bf y}^{\prime\prime}_{n}$ and ${\bf y}^{\prime\prime\prime}_{n}$ in the AMR boundary ghost zones on the finer grid, and then apply Eqs. (16)-(19) to compute ${\bf Y}_{i}$ in the AMR ghost region.
Since the finer grid requires two steps to synchronize with the coarser grid, we apply Eqs. (16)-(19) twice to interpolate the boundary values $Y_{i}$ on the finer grid at $t_{n}$ ($\theta=0$) and $t_{n}+h/2$ ($\theta=1/2$).
@@
    Perform a full update on the coarse level using the physical boundary condition to compute ${\bf Y}^{(0)}_{1}$, …, ${\bf Y}^{(0)}_{4}$ at the boundary. Store ${\bf k}^{(0)}_{1}$, …, ${\bf k}^{(0)}_{4}$.
@@
    On the next finer level $\ell$ prolongate in space ${\bf k}^{(\ell-1)}_{1}$, …, ${\bf k}^{(\ell-1)}_{4}$ and ${\bf y}_{n}^{(\ell-1)}$ to mesh refinement boundary.
@@
    Perform a full update on level $\ell$. Use (11)-(19) with $\theta=0$ to compute ${\bf Y}^{(\ell)}_{1}$, …, ${\bf Y}^{(\ell)}_{4}$ at the mesh refinement boundary. Use (11) with $\theta=1/2$ to compute ${\bf y}^{(\ell)}$ at mesh refinement boundary when implementing (6). Store ${\bf k}^{(\ell)}_{1}$, …, ${\bf k}^{(\ell)}_{4}$.
@@
    Perform a full update on level $\ell$. Use (11)-(19) with $\theta=1/2$ to compute ${\bf Y}^{(\ell)}_{1}$, …, ${\bf Y}^{(\ell)}_{4}$ at the mesh refinement boundary. Use (11) with $\theta=1$ to compute $y^{\ell}$ at mesh refinement boundary when implementing (6). Store ${\bf k}^{(\ell)}_{1}$, …, ${\bf k}^{(\ell)}_{4}$.
@@
As illustrated in Fig. 1, Carpet employs buffer zones to evolve fine grid points near AMR boundaries, requiring a prolongation width of four ghost zones for RK4 integration. In contrast, CarpetX utilizes the dense output method, which necessitates only a single ghost-zone-width prolongation. Carpet performs two prolongations of the state vector for every coarse time step, each involving $\sim$ 4 times as many points, while CarpetX performs prolongation once per coarse time step but for 5 times as many variables (the ${\bf k}_{1},\ldots,{\bf k}_{4}$ and the state vector). The net effect is CarpetX prolongates 5/8 as many points. In addition, because Carpet treats three of the buffer zones as interior points, the RHS calculation is performed over more points (slowing the calculation). The net effect is that CarpetX is faster and scales better, as seen in Sec. IV.3.
previous timefirst substepsecond substepthird substepfinal time
Figure 1: Integration schemes at AMR boundaries: (left) Carpet’s buffer zone and (right) CarpetX’s ghost zone treatment. The vertical dashed line separates interior points from AMR boundary zones. Left (Carpet): Empty circles ($\circ$) represent valid buffer zone points, while crosses ($\times$) represent points with invalid data during time integration. Each empty circle spans one ghost zone width, with each RK4 step requiring four ghost-zone-width prolongations. Right (CarpetX): Instead of excluding points, we prolongate the RK4 vectors ($\bf{y}_{n},\bf{k}_{1},...,\bf{k}_{4}$) across one ghost-zone width from the coarser level, then populate all ghost points at intermediate substeps and the final time using dense output method.
@@ III.2 Initial Data
III.3 Numerical Implementation
To stabilize numerical evolution, we incorporate KO dissipation  [34] into the system. This is implemented by modifying the time derivative of the evolved variables as follows:
$\partial_{t}u\rightarrow\textsc{RHS}+Qu$ (38)
where $Qu$ represents the dissipation term, given by:
@@ IV.1 Scalar Wave
where $f(v)\equiv A\exp(-v^{2}/(2\sigma^{2}))$, with $A=1.0$, $\sigma=0.25/\sqrt{2}$. The scalar wave equation was evolved at three different resolutions: $\Delta x=h_{1},h_{1}/1.2,h_{1}/1.44$, where $h_{1}=0.032$. The computational domain utilized a two-level refinement grid, with refinement boundaries at $r=4.0$, and $1.0$, respectively.
@@
Figure 2: Errors in the solution of a Gaussian scaler wave. The plots show errors in low- ($h_{1}=0.032$), medium- ($h_{2}=h_{1}/1.2$), and high-resolution ($h_{3}=h_{1}/1.44$) numerical solutions for both uniform time step and subcycling cases. The dashed lines represent the scaled medium- and high-resolution errors assuming fourth-order convergence, and they align closely with the low-resolution solution.
@@ IV.2.1 Convergence tests for Z4cow
Figure 3: Convergence test of a $q=1$, nonspinning binary using Z4cow. The top panel shows the low- ($h_{1}=3.3M$), medium- ($h_{2}=h_{1}/1.2$), and high-resolution ($h_{3}=h_{1}/1.44$) waveforms, which are visually indistinguishable, demonstrating excellent agreement across the resolutions. The middle and bottom panels display the waveform differences: the low- vs. high-resolution difference (solid black), the mid- vs. high-resolution difference (solid orange), and rescaled mid- vs. high-resolution difference (dashed blue), assuming fourth-order convergence.
In Fig.3, the top panel illustrates the evolution of the real part of the dominant $l=2$, $m=2$ mode of the Weyl scalar $\Psi_{4}$, extracted at a radius of $r=60M$, which encodes outgoing gravitational radiation through $\ddot{h}=\Psi_{4}$, where $h$ is the gravitational wave strain. The evolution is shown over the time interval $t=300M$ to $t=600M$ for three different resolutions. Here, $\Psi_{4}$ is calculated using the WeylScal4 [37] and Multipole thorns in the SpacetimeX repository. At this scale, the results are visually indistinguishable, indicating excellent agreement across the resolutions.
The middle and bottom panels display the differences between the low- and high-resolution results, as well as the medium- and high-resolution results. Additionally, the difference between the medium- and high-resolution results is rescaled under the assumption of fourth-order convergence. Numerical oscillations are observed around $t=70M$, $t=200M$, and $t=300M$, which are attributed to numerical errors that do not follow fourth-order convergence (middle panel). However, once the gravitational wave signal becomes stronger and dominates, starting around $t=500M$ , the results exhibit clear fourth-order convergence (bottom panel). This behavior confirms the robustness and accuracy of the implementation in capturing the dominant physical features of the system.
@@ IV.2.2 Comparison with LazEv
Figure 4: A comparison of the Z4cow and LazEv waveforms. The top panel shows the amplitudes of the waveforms from the two codes. The next panel down shows the relative differences in the amplitudes between the two codes. The third panel show the waveform phases for the two codes. Finally, the bottom panel show the phase differences between the two codes.
@@ IV.2.3 Irreducible mass
Figure 5: Irreducible mass of the inspiraling binary black holes (top) and the merged black hole (bottom), where $h_{1}=3.3M$, $h_{2}=h_{1}/1.2$, $h_{3}=h_{1}/1.44$. As the resolution increases, mass conservation improves.
@@ IV.2.4 Constraint violations
Figure 6: Hamiltonian and Momentum constraints for three different resolution cases, where $h_{1}=3.3M$, $h_{2}=h_{1}/1.2$, $h_{3}=h_{1}/1.44$. The dashed lines represent the scaled medium- and high-resolution constraints, assuming fourth-order convergence. Fourth-order convergence is observed, except around $t=200M$ and $t=400M$.
@@ IV.3 Scaling Test
$224$ $5.88\times 10^{7}$ 1.00 $224$ $1.14\times 10^{8}$ 1.00 $4$ $1.82\times 10^{9}$ 1.00
$448$ $1.07\times 10^{8}$ 0.91 $448$ $1.96\times 10^{8}$ 0.86 $8$ $2.60\times 10^{9}$ 0.71
$896$ $1.61\times 10^{8}$ 0.68 $896$ $3.26\times 10^{8}$ 0.71 $16$ $3.48\times 10^{9}$ 0.48
$1792$ $2.00\times 10^{8}$ 0.43 $1792$ $5.40\times 10^{8}$ 0.59 $32$ $4.03\times 10^{9}$ 0.28
$3584$ $1.69\times 10^{8}$ 0.18 $3584$ $8.90\times 10^{8}$ 0.49 $64$ $3.97\times 10^{9}$ 0.14
@@
$224$ $1.44\times 10^{7}$ 1.00 $224$ $2.85\times 10^{7}$ 1.00 $4$ $4.55\times 10^{8}$ 1.00
$448$ $1.41\times 10^{7}$ 0.98 $448$ $2.64\times 10^{7}$ 0.92 $8$ $4.26\times 10^{8}$ 0.94
$896$ $1.41\times 10^{7}$ 0.98 $896$ $2.49\times 10^{7}$ 0.87 $16$ $3.96\times 10^{8}$ 0.87
$1792$ $1.30\times 10^{7}$ 0.90 $1792$ $2.48\times 10^{7}$ 0.87 $32$ $3.94\times 10^{8}$ 0.87
$3584$ $1.19\times 10^{7}$ 0.83 $3584$ $2.45\times 10^{7}$ 0.86 $64$ $3.70\times 10^{8}$ 0.81
@@ Acknowledgements.
The authors would like to thank Manuela Campanelli, Michail Chabanov, Jay V. Kalinani, Carlos Lousto and Weiqun Zhang for valuable discussions. The authors gratefully acknowledge the National Science Foundation (NSF) for financial support from grants OAC-2004044, OAC-2004157, OAC-2004879 PHY-2110338, OAC-2411068, PHY02409706, as well as the National Aeronautics and Space Administration (NASA) for financial support from TCAN Grant No. 80NSSC24K0100. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. RH acknowledges additional support by NSF via grants OAC-2103680, OAC-2310548, and OAC-2005572. ES and RH acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). ES et RH remercie le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) de son soutien. This research used resources from the Texas Advanced Computing Center’s (TACC) Frontera and Vista supercomputer allocations (award PHY20010). Additional resources were provided by the BlueSky, Green Prairies, and Lagoon clusters of the Rochester Institute of Technology (RIT) acquired with NSF grants PHY-2018420, PHY-0722703, PHY-1229173, and PHY-1726215. We thank LSU HPC and CCT IT Services for making the mike compute resource available.